# RsGCN: Rescaling Enhances Generalization of GCNs for Solving Scalable Traveling Salesman Problems

### v3

@@ Abstract
Neural traveling salesman problem (TSP) solvers face two critical challenges: poor generalization for scalable TSPs and high training costs. To address these challenges, we propose a new Rescaling Graph Convolutional Network (RsGCN). Focusing on the scale-dependent features (i.e., features varied with problem scales) related to nodes and edges that influence the sensitivity of GCNs to the problem scales, a Rescaling Mechanism in RsGCN enhances the generalization capability by (1) rescaling adjacent nodes to construct a subgraph with a uniform number of adjacent nodes for each node across various scales of TSPs, which stabilizes the graph message aggregation; (2) rescaling subgraph edges to adjust the lengths of subgraph edges to the same magnitude, which maintains numerical consistency. In addition, an efficient training strategy with a mixed-scale dataset and bidirectional loss is used in RsGCN. To fully exploit the heatmaps generated by RsGCN, we design an efficient post-search algorithm termed Re2Opt, in which a reconstruction process based on adaptive weight is incorporated to help avoid local optima. Based on a combined architecture of RsGCN and Re2Opt, our solver achieves remarkable generalization and low training cost: with only 3 epochs of training on the mixed-scale dataset containing instances with up to 100 nodes, it can be generalized successfully to 10K-node instances without any fine-tuning. Extensive experiments demonstrate our state-of-the-art performance across uniform distribution instances of 9 different scales from 20 to 10K nodes and 78 real-world instances from TSPLIB, while requiring the fewest learnable parameters and training epochs among neural competitors.
@@ 1 Introduction
The traveling salesman problem (TSP), as a typical combinatorial optimization problem, has been extensively studied in the literature. Although classical solvers like Concorde [1] and LKH [7] achieve strong performance across various problem scales, their designs highly rely on expert-crafted heuristics. The development of deep learning has spawned numerous neural TSP solvers with various learning paradigms (e.g., Supervised Learning (SL), Reinforcement Learning (RL)) and architectural foundations (e.g., Graph Convolutional Networks (GCNs) [6, 9, 12, 18, 19, 27], Transformers [5, 15, 16, 17, 22, 24, 33], and Diffusion Models [18, 19, 27]). Herein, a typical framework is using neural networks to generate heatmaps and then incorporate post-search algorithms to obtain solutions with the guidance of heatmaps [6, 18, 19, 27]. Monte Carlo Tree Search [6], owing to its promising search efficiency, is frequently employed in the post-search guided by heatmaps.
@@
Although prior work has shown neural networks’ substantial potential in solving TSPs, it also reveals critical limitations in solution generalization [11]. Specifically, pre-trained models cannot perform well for new problem scales, and thus further fine-tuning with new-scale instances is necessary to transfer pre-trained models to the new problem scales. Existing approaches to enhance generalization include: (1) enhancing solution diversity by generating multiple heatmaps [18, 19, 27] or designing specific post-search [5, 6, 8, 9], and (2) decomposing large-scale problems into small-scale sub-problems for partitioned solving [6, 17, 22, 24, 33]. Although these approaches achieve modest improvements, they suffer from prohibitive computational overhead. Regarding (1), generating additional heatmaps is computationally intensive for GPUs, significantly diminishing the practicality. Furthermore, existing post-search algorithms exhibit insufficient capability in narrowing down the search space and thus result in limited search efficiency. As for (2), the decomposition of large-scale problems leads to a partial loss of global perspective and thus weakens the overall optimization effectiveness, while the design of sub-problems partitioning and recombination introduces significant complexity.
@@
Based on our research analysis, since neural networks are sensitive to numerical values, the scale-dependent features related to nodes and edges in scalable TSPs are essential factors impairing generalization. In this paper, we propose a new GCN termed RsGCN, which incorporates a Rescaling Mechanism to adaptively rescale the scale-dependent features. As shown in Figure 1, the Rescaling Mechanism (1) rescales adjacent nodes to stabilize the graph message aggregation, which constructs a subgraph for each node by k-Nearest Neighbor Selection to ensure each node has a uniform number of adjacent nodes regardless of instance scale; (2) rescales subgraph edges to maintain numerical consistency, which adjusts the length of edges in subgraphs to the same magnitude by Uniform Unit Square Projection. With the Rescaling Mechanism, RsGCN can learn universal patterns by training on only small-scale instances. Thus, RsGCN can be effortlessly generalized to large-scale TSPs and generate high-quality heatmaps to guide post-search for scalable TSPs. In addition, rethinking the limitations of existing post-search algorithms, we design a new post-search algorithm termed Reconstruction-2Opt (Re2Opt), which utilizes 2Opt as the basic optimizer and performs reconstruction based on adaptive weights to robustly escape from local optima.
According to extensive experimental studies, the combined architecture of RsGCN and Re2Opt shown in Figure 1 achieves state-of-the-art performance on both uniform distribution and real-world instances. Our contributions are summarized as follows:
  1. 1.
    Rescaling Mechanism for Enhancing Generalization: We propose a Rescaling Mechanism to rescale the number of adjacent nodes and the lengths of subgraph edges, enhancing GCN into RsGCN to significantly improve generalization capability in solving scalable TSPs. With the Rescaling Mechanism, RsGCN requires significantly fewer learnable parameters and training epochs compared with other neural solvers and only uses a mixed-scale dataset containing instances with up to 100 nodes for training, while the obtained model can be effectively generalized to large-scale instances (e.g., 10K-node instances) without any fine-tuning.
  2. 2.
    Efficient Post-Search for Leveraging Heatmaps: To sufficiently utilize the informative heatmaps generated by RsGCN, we design a new Re2Opt algorithm for post-search, which incorporates reconstruction to help escape from local optima.
  3. 3.
    Ordered Heatmaps for Visual Analysis: We introduce ordered heatmaps to facilitate the visual analysis of the heatmaps’ quality and help validate the effectiveness of the Rescaling Mechanism.
Figure 1: The architecture of RsGCN with heatmap-guided post-search Re2Opt.
@@ 2 Related Work
Based on the differences in the way of constructing solutions, existing neural TSP solvers can be divided into two categories: rule-based constructive solvers and learning-based constructive solvers.
2.1 Rule-Based Constructive (RC) Solvers
In RC solvers [12, 6, 9, 27, 18, 19], the trained neural networks output an informative heatmap that indicates promising connections between nodes. Subsequently, the heatmap is used to guide the post-search algorithm to construct the TSP tour. In other words, neural networks are only involved in the generation of the heatmap and do not directly obtain valid solutions for TSPs. The encoders of RC solvers are typically based on GCN, while MLP [12, 6, 9] or Diffusion [27, 18, 19] are employed as decoders.
In the post-search with heatmaps, simple algorithms (e.g., greedy search and beam search) are used to generate TSP tours as initial states [12, 9, 27, 19] and 2Opt [4] algorithm is used to further optimize the generated TSP tours [27, 19]. In addition, diffusion-based models generate multiple heatmaps with different random seeds and then conduct greedy search to obtain multiple solutions. To perform search with a broader range based on the heatmap, Monte Carlo Tree Search (MCTS) [6] is proposed based on a Markov decision process to perform k-Opt operations. In general, both high-quality heatmaps and efficient post-search algorithms are crucial in RC solvers.
2.2 Learning-Based Constructive (LC) Solvers
LC solvers [15, 16, 5, 24, 22, 33, 17] take the current solution state as the input of neural networks and output the complete TSP tours. Such a solving process usually involves divide-and-conquer and inference within multiple iterations, which are mainly learned by neural networks. For example, H-TSP [24] employs a hierarchical approach with lower-level and upper-level models, in which the lower-level model solves sub-problems while the upper-level model combines the solutions of sub-problems into a complete TSP tour. DRHG [17] breaks edges within a cluster and constructs a sub-problem based on a hyper-graph structure before optimization. Transformer-based models with autoregressive construction have been effectively applied in LC solvers. Note that although LC solvers also use some simple rule-based search algorithms to construct solutions, most of the computational overhead occurs in the autoregressive and iterative process of neural networks.
2.3 Rethinking Complexity and Generalization
The heatmaps obtained by RC solvers do not show sufficient generalization capability in scalable TSPs and thus need to be fine-tuned with multiple new-scale instances to adapt to scalable TSPs, which increases the training cost. To enhance solution diversity, the diffusion-based decoder is used [27, 18, 19], which generates multiple heatmaps with multiple random seeds. However, the multi-step denoising process in the diffusion-based decoder is computationally intensive during training and inference. In addition, although some post-search algorithms such as MCTS achieve promising performance, their hyperparameters should be tuned for different scales of TSPs, which relies on expert knowledge and weakens the generalization capability. A recent study [31] has also questioned the role of heatmaps in guiding MCTS because MCTS can perform better without heatmaps generated by neural networks.
LC solvers typically conduct divide-and-conquer to decompose a large-scale TSP into several small-scale sub-problems [24, 22, 33, 17], and thus they have a certain capability of generalization for scalable TSPs. However, the optimization of each sub-problem loses the global perspective and may easily obtain a local optimum. If no problem decomposition is conducted or the scale of sub-problems is enlarged, which can enhance the global search capability, the computational complexity of LC solvers will extensively increase and result in unacceptable solution time overhead. In addition, the decomposition of sub-problems also relies on expert knowledge. In [24, 33], a hierarchical problem-solving framework is employed, which combines multiple models trained on different scales of TSP datasets to enhance the generalization capability. However, such a framework is complex and requires expensive computational costs.
To enhance the generalization capability while reducing the complexity and computational cost, we incorporate a Rescaling Mechanism into GCN and propose a new RsGCN, which learns universal patterns across various problem scales by rescaling the scale-dependent features related to nodes and edges in TSPs. With only 3 training epochs on a mixed-scale dataset (including instances with 20, 30, 50, and 100 nodes), RsGCN can be successfully generalized to 10K-node instances without any fine-tuning. An experimental investigation is also conducted to validate the efficient guidance provided by the heatmaps of RsGCN. In addition, a new post-search algorithm Re2Opt is proposed, which does not require dedicated hyperparameter tuning for different scales of TSPs.
@@ 3 Preliminaries
In TSPs, nodes are located in a 2-dimensional Euclidean space with undirected edges between nodes. For an $n$-node instance, $\mathcal{X}_{n}=\{x_{1},x_{2},\dots,x_{n}\}$ denotes its set of node coordinates, where $x_{i}=(a_{i},b_{i})$. A solution for TSPs, i.e., a TSP tour, is a closed loop that traverses all nodes exactly once and is typically encoded by a permutation $\Pi_{n}=\{\pi_{1},\pi_{2},\dots,\pi_{n}\}$. The length of the tour $\Pi_{n}$, denoted by $L(\Pi_{n})$, is calculated as follows, and the optimization objective of TSP is finding a permutation with the minimum tour length.
$L(\Pi_{n})=\sum_{i=1}^{n-1}(\|x_{\pi_{i}}-x_{\pi_{i+1}}\|^{2})+\|x_{\pi_{n}}-x_{\pi_{1}}\|^{2}$ (1)
@@
GCNs output a heatmap $\boldsymbol{H}\in\mathbb{R}^{n\times n}$ with input of an $n$-node instance, where $H_{i,j}$ represents the heat between $x_{j}$ and $x_{i}$, i.e., the probability that $x_{j}$ is adjacent to $x_{i}$ in the generated TSP tour. To facilitate observation and analysis, we define ordered heatmaps that are sorted according to the optimal tour. Details of ordered heatmaps are shown in Appendix A.
@@ 4 Methodology
4.1 Rescaling Mechanism for GCN (RsGCN)
GCNs have powerful feature representation capability to generate high-quality heatmaps for fixed-scale TSPs. However, previous studies [11, 27] have shown that GCNs exhibit limited generalization capability for scalable TSPs. We find that the primary issue stems from GCNs’ sensitivity to scale-dependent features, including (1) the number of adjacent nodes, and (2) the measurement of distances from any given node to its adjacent nodes. Specifically, the scale-dependent features vary in different scales of TSPs and thus weaken the generalization capability. To enhance the generalization capability for scalable TSPs, the proposed RsGCN incorporates the Rescaling Mechanism to rescale the scale-dependent features and help learn universal patterns across TSPs with different scales, which consists of rescaling adjacent nodes and rescaling subgraph edges as follows.
4.1.1 Rescaling Adjacent Nodes
@@
To address this issue, we employ k-Nearest Neighbor Selection to rescale the number of adjacent nodes as $k_{1}$ for each node, thereby each node forms a subgraph including itself and its top $k_{1}$ nearest neighbors. $k_{1}$ is fixed and thus the nodes in TSPs with various scales have subgraphs with a uniform number of adjacent nodes, helping learn universal patterns from instances with different scales. In addition, it can also achieve graph sparsification [20] that enhances GCNs by retaining promising candidate nodes and stabilizing message aggregation.
@@
Specifically, for an $n$-node instance with node set $\mathcal{X}_{n}=\{x_{1},x_{2},\dots,x_{n}\}$, the subgraph set is denoted by $\boldsymbol{\mathcal{N}}(\mathcal{X}_{n}|k_{1})=\{\mathcal{N}_{1}^{k_{1}},\mathcal{N}_{2}^{k_{1}},\dots,\mathcal{N}_{n}^{k_{1}}\}$, where $\mathcal{N}_{i}^{k_{1}}=\{x_{i}^{1},x_{i}^{2},\dots,x_{i}^{k_{1}}\}$ contains the $k_{1}$-nearest neighbors of $x_{i}$ (including $x_{i}$ itself) and represents $x_{i}$’s subgraph. $x_{i}^{j}=(a_{i}^{j},b_{i}^{j})\in\mathcal{X}_{n}$ denotes the coordinate of $x_{i}$’s $j$-th nearest neighbor. Note that after rescaling the adjacent nodes, the range of message aggregation is narrowed down and the complexity of message aggregation in GCNs is reduced from $O(n^{2})$ to $O(nk_{1})$.
4.1.2 Rescaling Subgraph Edges
In existing research, the coordinates of nodes in TSPs are typically normalized in a unit square space $[0,1]^{2}$ to uniform the range of node distribution and thus can facilitate the training of neural networks. However, such a normalization also influences the edge features in TSPs, which is neglected. In detail, with the same range of space, the distribution of nodes in an instance with more nodes is denser, which makes the distances between nodes smaller on average. Therefore, the edge features about distances between nodes vary in scalable TSPs, which limits the generalization capability of neural TSP solvers.
To deal with the above issue, the Uniform Unit Square Projection is proposed and incorporated into the Rescaling Mechanism. For a subgraph of a node $x_{i}$, the overall principle is to rescale the coordinates of nodes in $x_{i}$’s subgraph and obtain a transformed subgraph in which the coverage of $x_{i}$ and its neighbors is maximized in the unit square space.
@@ 4.1.2 Rescaling Subgraph Edges
In detail, we first acquire the maximum and minimum coordinate values of nodes on the two axes from $\mathcal{N}_{i}^{k_{1}}$, denoted by $a_{i}^{\max}$, $a_{i}^{\min}$, $b_{i}^{\max}$, and $b_{i}^{\min}$, respectively. Then, we compute the scaling factor $\mu_{i}$ as $\mu_{i}=\frac{1}{\max(A_{i},B_{i})}$, where $A_{i}=a_{i}^{\max}-a_{i}^{\min}$ and $B_{i}=b_{i}^{\max}-b_{i}^{\min}$. Based on the above calculations, we rescale each node coordinate $x_{i}^{j}=(a_{i}^{j},b_{i}^{j})$ in $\mathcal{N}_{i}^{k_{1}}$ to obtain the rescaled coordinate $r(x_{i}^{j})$ according to Equation 2 and finally project $\mathcal{N}_{i}^{k_{1}}$ into $R(\mathcal{N}_{i}^{k_{1}})$ according to Equation 3. Herein, $\mu_{i}$ is used to maximize the coverage of node distribution in the unit square space while ensuring that the rescaled coordinate does not exceed the boundary of the unit square space.
$r(x_{i}^{j})=r[(a_{i}^{j},b_{i}^{j})]=((a_{i}^{j}-a_{i}^{\min})\times\mu_{i},\ (b_{i}^{j}-b_{i}^{\min})\times\mu_{i}),\ \forall j\in\{1,2,\dots,k_{1}\}$ (2)
$R(\mathcal{N}_{i}^{k_{1}})=\{r(x_{i}^{1}),r(x_{i}^{2}),\dots,r(x_{i}^{k_{1}})\},\ \forall i\in\{1,2,\dots,n\}$ (3)
After rescaling subgraph edges, we calculate the Euclidean distances from $x_{i}$ to its $k_{1}$-nearest neighbors based on $R(\mathcal{N}_{i}^{k_{1}}),\forall i\in\{1,2,\dots,n\}$. Consequently, the edge lengths in each subgraph from TSP instances with different scales are rescaled to the same magnitude, reducing the influence of distance measurement on the generalization capability in dealing with scalable TSPs. For intuitive comprehension, an example of the Uniform Unit Square Projection is shown in Figure 2. The leftmost figure represents the original distribution of the nodes in a subgraph, and the red rectangle denotes the subgraph frame constructed based on the marginal nodes of the subgraph. After the two steps illustrated in Figure 2, the rescaled subgraph is obtained and is shown in the rightmost figure. Note that each arrow represents a step, with the specific operation displayed above and the related variables displayed below the arrow. In addition, the effect of the Rescaling Mechanism is visualized and discussed in more detail in Appendix H.

### v4

# RsGCN: Subgraph-Based Rescaling Enhances Generalization of GCNs for Solving Traveling Salesman Problems

@@ Abstract
GCN-based traveling salesman problem (TSP) solvers face two critical challenges: poor cross-scale generalization for TSPs and high training costs. To address these challenges, we propose a Subgraph-Based Rescaling Graph Convolutional Network (RsGCN). Focusing on the scale-dependent features (i.e., features varied with problem scales) related to nodes and edges, we design the subgraph-based rescaling to normalize edge lengths of subgraphs. Under a unified subgraph perspective, RsGCN can efficiently learn scale-generalizable representations from small-scale TSPs at low cost. To exploit and assess the heatmaps generated by RsGCN, we design a Reconstruction-Based Search (RBS), in which a reconstruction process based on adaptive weight is incorporated to help avoid local optima. Based on a combined architecture of RsGCN and RBS, our solver achieves remarkable generalization and low training cost: with only 3 epochs of training on a mixed-scale dataset containing instances with up to 100 nodes, it can be generalized successfully to 10K-node instances without any fine-tuning. Extensive experiments demonstrate our advanced performance across uniform-distribution instances of 9 different scales from 20 to 10K nodes and 78 real-world instances from TSPLIB, while requiring the fewest learnable parameters and training epochs among neural competitors.
@@ 1 Introduction
The traveling salesman problem (TSP), as a typical combinatorial optimization problem, has been extensively studied in the literature. Although classical solvers like Concorde (Applegate et al., 2006) and LKH (Helsgaun, 2000) achieve strong performance across various problem scales, their designs highly rely on expert-crafted heuristics. The development of deep learning has spawned numerous neural TSP solvers with various learning paradigms (e.g., Supervised Learning, Reinforcement Learning) and architectural foundations (e.g., Graph Convolutional Networks (GCNs) (Joshi et al., 2019; Fu et al., 2021; Hudson et al., 2022; Sun & Yang, 2023; Li et al., 2023; 2024) and Transformers (Kool et al., 2019; Kwon et al., 2020; Drakulic et al., 2023; Luo et al., 2023; Pan et al., 2023; Ye et al., 2024; Li et al., 2025)). Herein, a typical framework is using neural networks to generate heatmaps and then incorporate post-search algorithms to obtain solutions with the guidance of heatmaps. LKH-3 (Helsgaun, 2017) and Monte Carlo Tree Search (Fu et al., 2021) are commonly employed in the post-search guided by heatmaps.
@@
Although prior work has shown substantial potential of GCNs in solving TSPs, it also reveals critical limitations in solution generalization (Joshi et al., 2021). Specifically, pre-trained models cannot perform well for new problem scales, and thus further fine-tuning with new-scale instances is necessary to transfer pre-trained models to the new problem scales. Existing approaches to enhance generalization include: (1) enhancing solution diversity by generating multiple heatmaps (Sun & Yang, 2023; Li et al., 2023; 2024), and (2) decomposing large-scale problems into small-scale sub-problems for partitioned solving (Fu et al., 2021; Luo et al., 2023; Pan et al., 2023; Ye et al., 2024; Li et al., 2025). Although these approaches achieve modest improvements, they suffer from prohibitive computational overhead. Regarding (1), generating additional heatmaps is computationally intensive for GPUs, significantly diminishing the practicality. As for (2), the decomposition of large-scale problems leads to a partial loss of global perspective and thus weakens the overall optimization effectiveness, while the design of sub-problems partitioning and recombination introduces additional complexity.
@@
Based on our research analysis, since GCNs are sensitive to numerical values, the scale-dependent features related to nodes and edges in TSPs are essential factors impairing generalization. In this paper, we propose a new GCN termed RsGCN, which incorporates the subgraph-based rescaling to adaptively rescale subgraph edges. As shown in Figure 1, the subgraph-based rescaling first (1) construct subgraphs for each node by k-Nearest Neighbor (k-NN) selection to ensure each node has a uniform number of neighbors regardless of instance scale; then (2) rescales subgraph edges by Uniform Unit Square Projection to maintain consistency of numerical distributions. With the subgraph-based rescaling, RsGCN can learn universal patterns by training on only small-scale instances. Thus, RsGCN can be effortlessly generalized to large-scale TSPs and generate high-quality heatmaps to guide post-search more efficiently. In addition, rethinking the limitations of existing post-search algorithms, we design a new post-search algorithm termed Reconstruction-Based Search (RBS), which utilizes 2-Opt as the basic optimizer and performs reconstruction based on adaptive weights to robustly escape from local optima. According to extensive experimental studies, the integrated framework of RsGCN and RBS shown in Figure 1 achieves state-of-the-art performance on both uniform-distribution and real-world instances among neural competitors.
Figure 1: The integrated framework of RsGCN and RBS.
@@ 2 Related Work
According to the model backbones, neural TSP solvers can be broadly categorized into Transformer-based and GCN-based methods. In this work, we focus on the generalization and cost of GCN-based methods.
GCN-based methods typically employ an edge-aware variant of GCN, namely Gated GCN (Bresson & Laurent, 2018). For GCN-based methods (Joshi et al., 2019; Fu et al., 2021; Xin et al., 2021; Hudson et al., 2022; Sun & Yang, 2023; Li et al., 2024), the trained GCN outputs informative heatmaps that indicate promising candidates. Subsequently, heatmaps are used to guide the post-search algorithm to construct the TSP tour. Much of the related research has centered on improving the generalization performance of GCNs. Joshi et al. (2021); Xin et al. (2021); Lischka et al. (2024) demonstrated that graph sparsification based on k-NN can improve the generalization of GCNs to some extent. Fu et al. (2021) proposed decomposing large-scale problems into multiple smaller sub-problems for separate solving, and then merging them using Monte Carlo Tree Search (MCTS). Moreover, Sun & Yang (2023); Li et al. (2024) employed diffusion models as decoders to generate diverse heatmaps, thereby enhancing solution diversity.
However, the decomposition-merging paradigm lacks a global perspective, leading to degraded solution quality. Incorporating Diffusion models as decoders introduces additional computational overhead while offering no substantive improvement to the graph encoder. Overall, prior work has overlooked the high sensitivity of edge-aware GCNs to edge weights, which is the essential factor limiting their generalization. Notably, a Transformer-based method, INViT (Fang et al., 2024), introduces invariant nested views to normalize node distributions, thereby enhancing the generalization for cross-distribution TSP. Building on prior studies and analyses, we propose a subgraph-based edge normalization method and design RsGCN, which incorporates a global perspective and strong cross-scale generalization. We empirically demonstrate that RsGCN can effectively generalize from small-scale to large-scale TSPs with very few parameters and low training cost.
@@ 3 Preliminaries
In symmetric TSPs, nodes are located in a 2-dimensional Euclidean space with undirected edges. For an $n$-node instance, $\mathcal{X}_{n}=\{x_{1},x_{2},\dots,x_{n}\}$ denotes its set of node coordinates, where $x_{i}=(a_{i},b_{i})$. A solution for TSPs, i.e., a TSP tour, is a closed loop that traverses all nodes exactly once and is typically encoded by a permutation $\Pi_{n}=\{\pi_{1},\pi_{2},\dots,\pi_{n}\}$. The length of the tour $\Pi_{n}$, denoted by $L(\Pi_{n})$, is calculated as in Equation 1, where $\|\cdot\|$ denotes the Euclidean norm. Finally, the optimization objective is to find the permutation that minimizes the tour length.
$L(\Pi_{n})=\sum_{i=1}^{n-1}(\|x_{\pi_{i}}-x_{\pi_{i+1}}\|)+\|x_{\pi_{n}}-x_{\pi_{1}}\|.$ (1)
@@
GCNs output a heatmap $\boldsymbol{H}\in\mathbb{R}^{n\times n}$ with input of an $n$-node instance, where $\boldsymbol{H}_{i,j}$ represents the heat of the edge $x_{j}\rightarrow x_{i}$, i.e., the probability that $x_{j}$ is the successor of $x_{i}$ in the predicted TSP tour. Indicative heatmaps can guide the post-search process, narrowing the search space and thereby improving solution quality.
@@ 4 Methodology
4.1 Subgraph-Based Rescaling for GCN – RsGCN
GCNs have powerful feature representation capability to generate high-quality heatmaps for fixed-scale TSPs. However, previous studies (Joshi et al., 2021; Sun & Yang, 2023) have shown that GCNs exhibit limited cross-scale generalization capability. Specifically, scale-dependent features vary in different scales of TSPs and thus weaken the generalization capability. To learn universal patterns and enhance cross-scale generalization, our RsGCN incorporates the subgraph-based rescaling to rescale subgraph edges, which consists of two steps as follows.
4.1.1 Construct Subgraphs
@@
To address this issue, we employ k-NN selection to prune the number of adjacent nodes as $k_{1}$ for each node, thereby each node forms a subgraph including itself and its top $k_{1}$ nearest neighbors. $k_{1}$ is fixed and thus the nodes in TSPs with various scales have subgraphs with a uniform number of adjacent nodes, helping learn universal patterns from instances with different scales. In addition, it can also achieve graph sparsification (Joshi et al., 2021; Xin et al., 2021; Lischka et al., 2024) that enhances GCNs by retaining promising candidates and stabilizing message aggregation.
@@
Specifically, for an $n$-node instance with node set $\mathcal{X}_{n}=\{x_{1},x_{2},\dots,x_{n}\}$, the subgraph set is denoted by $\boldsymbol{\mathcal{N}}(\mathcal{X}_{n}|k_{1})=\{\mathcal{N}_{1}^{k_{1}},\mathcal{N}_{2}^{k_{1}},\dots,\mathcal{N}_{n}^{k_{1}}\}$, where $\mathcal{N}_{i}^{k_{1}}=\{x_{i}^{1},x_{i}^{2},\dots,x_{i}^{k_{1}}\}$ contains the $k_{1}$-nearest neighbors of $x_{i}$ (including $x_{i}$ itself) and represents $x_{i}$’s subgraph. $x_{i}^{j}=(a_{i}^{j},b_{i}^{j})\in\mathcal{X}_{n}$ denotes the coordinate of $x_{i}$’s $j$-th nearest neighbor.
4.1.2 Rescale Subgraph Edges
In existing research, the coordinates of nodes in TSPs are typically normalized in a unit square space $0,1]^{2}$. However, with the same range of space, node distributions tend to become denser as the problem scale increases, which makes edge lengths smaller on average. Therefore, cross-scale differences in edge lengths can impair the cross-scale generalization of GCNs.
@@ 4.1.2 Rescale Subgraph Edges
To deal with the above issue, we introduce Uniform Unit Square Projection to rescaling subgraph edges. For a subgraph $\mathcal{N}_{i}^{k_{1}}$, the main idea is to project its node coordinates so that the coverage of $\mathcal{N}_{i}^{k_{1}}$ is maximized in the unit square space. In detail, we first acquire the maximum and minimum coordinate values on the two axes from $\mathcal{N}_{i}^{k_{1}}$, denoted by $a_{i}^{\max}$, $a_{i}^{\min}$, $b_{i}^{\max}$, and $b_{i}^{\min}$, respectively. Then, we compute the scaling factor $\mu_{i}$ and rescale subgraph edges according to Equations 2 and 3.
$\mu_{i}=\frac{1}{\max(a_{i}^{\max}-a_{i}^{\min},b_{i}^{\max}-b_{i}^{\min})}$ (2)
$\text{Rescale}(\|x_{i}-x_{i}^{j}\|)=\|x_{i}-x_{i}^{j}\|\cdot\mu_{i},\ \forall j\in\{1,2,\dots,k_{1}\}.$ (3)
After rescaling the edges of all subgraphs in $\boldsymbol{\mathcal{N}}(\mathcal{X}_{n}|k_{1})$, the cross-scale differences in edge lengths are eliminated. For intuitive comprehension, an example of the subgraph-based rescaling is shown in Figure 2. The leftmost figure shows the original node distribution, while the red rectangle denotes the subgraph frame. The rightmost figure presents the rescaled subgraph. In addition, the effect of the subgraph-based rescaling is visualized and discussed concretely in Appendix G.